OpenStudy (anonymous):

Which polynomial will definitely have nonreal zeros? Hint: Use Descartes's rule of signs. A. x^3-2x^2-x-2 B.X^3-2x^2+x-2 C.x^3-2x^2 D.X^4-8x^2+16

OpenStudy (anonymous):

@amoodarya

OpenStudy (amoodarya):

X^3-2x^2+x-2 =(x-2)(.....) x^3-2x^2=x^2(x-2) X^4-8x^2+16=(x^2-4)^2=(x-2)^2 *(x+2)^2

OpenStudy (anonymous):

what?

OpenStudy (amoodarya):

A,B has a one real root , and two unreal root(zero) but c, d have all real

OpenStudy (amoodarya):

one of them is correct ? do you type question (A,B) correct?

OpenStudy (anonymous):

I can only pick one. This is too hard

OpenStudy (amoodarya):

I think question is not correct or you type it wrong !

OpenStudy (amoodarya):

BOTH OF A,B HAS ONE real zero , and 2 unreal zero may be D is X^4+8x^2+16 instead of X^4-8x^2+16 so X^4+8x^2+16=(x^2+4)^2 and has no real zero

OpenStudy (anonymous):

well it wasnt B, must be A